School administrators wondered whether class size and grade achievement (in percent) were

related. A random sample of classes revealed the following data where the number of students, x,

affects the average grade, y. Given that ^y = 90.92 - 0.39x

No. of students: 15 10 8 20 18 6

Avg. grade (%): 85 90 82 80 84 92

a) Predict the average grade for the number of students is 20 and compute the residual. Is the

average grade average or below average for the number of students?

b) Predict the average grade for the number of students is 10 and compute the residual. Is the

average grade average or below average for the number of students?
Answer #1

a)

from given regression equation:

predicted value =90.92-0.39*20 =83.12

residual =actual-predicted value =80-83.12 =-3.12

since residual is negative ; below average for the number of students

b)

predicted value =90.92-0.39*10 =87.02

residual =actual-predicted value =90-87.02 =2.98

since residual is positive ; above average for the number of students